# Minimal cellular automaton model with heterogeneous cell sizes predicts epithelial colony growth

### v1

@@ Abstract
Understanding the principles of dynamic tissue organization requires an interdisciplinary effort where quantitative biological observations must be challenged by mathematical models that integrate existent insight on a theoretical level SchBru2023 ; BruArlFinRonRaeBro2021 . Cell-based mathematical models essentially consider a tissue as ’a society of cells’ Vir1855 where cells behave and interact with each other and the extracellular environment according to individual cellular characteristics summarizing the effects of sub-cellular processes, such giving rise to complex emergent processes at tissue scale. Existing cell-based model frameworks can be classified into three broad categories, (i) continuum descriptions of tissue movement which rarely describe individual cells explicitly AlaHatLowVoi2015 ; KuaGuaTanZha2023 ; Shr2005 ; BasRisJoaSasPro2009 ; HeiAleLaCZajKosCoh2020 ; KhaStr2021 , (ii) models that include an explicit description of the cell membrane, such as the cellular Potts model GraGla1992 ; LiLow2014 ; Vos2012 ; Kab2012 , vertex models NesGodWooJen2018 ; AltGanSal2017 ; FarRoeAigEatJue2007 ; KurBroZarBakFle2015 ; MorStuSabLauQueTorChaSuhRam2017 ; FleOstBakShv2014 or disk models MorStuSabLauQueTorChaSuhRam2017 ; SchMolYam2020 , and finally (iii) particle based-models or cellular automata where cells are represented as point-like particles BoeHatVosCavHerDeu2015 ; Vos2012 ; TalCavVosDeu2017 ; RosBoeLanVos2021 ; RehKliDeuVos2017 ; LanMogHueVos2022 ; HofVosRinBru2017 ; FraAlaBoeVosLan2022 ; SepPetCocGraSilHak2013 . Each class has its own merits and drawbacks: (i) Tissue scale descriptions allow to capture biophysical principles which govern large scale phenomena like tissue fluidity and deformations (elastic, plastic, viscoelastic) or jamming transitions YanBiCzaMerManMar2017 ; GniSchHal2019 . However, relating biophysical parameters of cell colonies to measurable single cell properties remains challenging and simulation of large collectives with single cell resolution can cause high computational cost AlaHatLowVoi2015 ; KuaGuaTanZha2023 . (ii) Models that describe cells as spatially extended objects, such as cellular Potts and vertex models, can incorporate heterogeneous cell sizes and shapes within the tissue and provide a related contact network for specifying intercellular interaction. However, the cellular behavior and interaction is often described indirectly via membrane adjustments controlled by free energy terms, the latter also involving purely technical parameters which have an impact on the emergent dynamics KurBakFle2017 ; Vos2012 ; BeaKirHenGraBru2023 . This fact together with the considerable computational effort for full exploration of the parameter space limits the calibration to concrete experimental data. (iii) Cellular automata and interacting particle systems are computationally more efficient while being mathematically easier to analyze LiePalJagDra2015 . They allow to describe cellular behavior and intercellular interaction in a rule-based manner based on clear mechanistic hypotheses and with biologically meaningful parameters OsbFlePitMaiGav2017 . A major drawback, however, is the idealization of point-like cells hampering the modeling of tissue dynamics where cell sizes and shapes are supposed to play an essential role AleTre2020 .
@@
Here, we consider monolayers of growing epithelial tissue viewed as a dynamic 2D arrangement of individual cells, which grow in size, divide and migrate in response to mechanical constraints and, potentially, interaction on cell-cell contact. In-vitro experiments on the development of epithelial tissue of Madin-Darby canine kidney cells from Puliafito et al. (2012) show two distinct growth regimes: At early times, in the confluent phase, cell density is low and spatially homogeneous and the colony grows exponentially in time. At longer times, in the postconfluent phase, cell density increases inside the colony and subexponential colony growth driven by more motile cells near the edge of the expanding colony is observed. The inhibition of cellular proliferation and migration in crowded environments as observed in the center of the growing monolayers is called contact inhibition of proliferation response and migration. It was found to be a not only a consequence of cell-cell contact but of mechanical constraints that cause successive cell divisions to reduce cell area. To further differentiate between the impact of biomechanical constraints and that of intercellular interaction on cell-cell contact, cell-based mathematical modeling was exploited. Puliafito et al. (2012) used a 1D vertex model to support their findings qualitatively. However, the model parameters and the observables of the emergent behavior were not systemically related to experimental measurements. Aland et al. (2015) could reproduce the observed cell culture topology and radial distribution function with a hybrid continuum-discrete model, however, the spatial scale of the culture had to be extrapolated to realistic culture sizes due to computational cost of the model. Jain et al. (2022) incorporated proliferation in the energetic description of a multiphase field model to predict the coordination numbers of cells in growing epithelial colonies corresponding to different hypotheses on contact inhibition and to confirm the typical, experimentally observed linear boundary growth of the colony radius. The role of cell polarity and cellular cycling between motile and dividing states for contact inhibition of proliferation was explored using a disk model SchMolYam2020 . However, the question whether mechanical constraints alone, without further intercellular interaction, can explain the observed tissue-scale dynamics remains still largely open.
@@
We study this question by proposing a minimal model, an extension of probabilistic cellular automata (CA) that allows describing heterogeneous cell sizes and growth of individual cells. In our model, cells located on the nodes of a 2D regular grid can grow, divide and migrate, see Fig. 1 for an illustration. They only interact by the exclusion princple, which means that a cellular action is prevented if there is insufficient space. We calibrate the model parameters, which quantify single cell dynamics, to previous in-vitro observations based on single cell measurements, and subsequently compare the emergent features of the calibrated CA model, such as culture size and cell density, to independent experiments. These features of CA model reproduce quantitatively the corresponding experimental results, implying that the exclusion principle, the only intercellular interaction incorporated in the model, paired with size dependent proliferation rates is sufficient to generate the observed contact inhibition. We thus demonstrate the effectiveness of the CA model for efficient quantitative predictions of tissue organization on the basis of rule-based cellular behavior with heterogeneous cell sizes. We discuss the potential to incorporate effective biomechanical interactions between cells in our CA model for future studies of contact inhibition.
@@
Figure 1: Illustration of the cellular automaton (CA) model. Cells (red and green) with individual sizes $A_{i}$ are positioned on a regular grid (gray $3\times 3$ squares represent a section of this grid), where each square of the grid can be occupied by several cells as long as their total size does not exceed the size of the square $\sum_{i}A_{i}\leq A_{\text{max}}$. Each cell $i$ can proliferate with rate $\gamma(A_{i})$, migrate with rate $\mu$, and grow logistically with rate $\alpha$ to a maximal size of $A_{\text{max}}$ (each action exemplarily sketched for the red highlighted cell in the central square). Cells interact via exclusion principle, i.e., migration and growth of individual cells is inhibited by lack of sufficient free space. Dynamics are numerically integrated using Markov-chain-Monte-Carlo method with time increment $\Delta t$. Note that the position of a cell is not resolved beyond its grid square.
@@
Figure 2: Exemplary simulation from the CA model of the growth of epithelial tissue in the confluent regime (1 d, 3.3 d, 5 d), as well as at (6.2 d, 7.3 d) and post-transition (8.3 d, 8.8 d, 9.8 d) to a stationary postconfluent state. For each time point, the local cell density is encoded in color. The growing culture expands outwards over time. Around day 6 the density increases drastically in the center of the culture, as the lack of free space inhibits migration and cell growth, reflecting the transition to a postconfluent state. At the same time, a radial density gradient is visible, implying that cells at the boundary of the culture are able to grow larger and drive further outwards expansion. Model parameters calibrated as written in the main text.
I CA model
@@
We present an efficient, agent-based dynamical model to simulate the development of two-dimensional tissue. The model is based on a probabilistic cellular automaton (CA) but additionally allows for size changes of individual cells during cell growth and division, see Fig. 1 for an illustration of the model and Fig. 2 for an exemplary simulation of epithelial tissue. In the CA, the two-dimensional space is discretized into a regular grid of $M^{2}$ squares each with size $A_{\text{max}}$, which can be occupied by cells. Several cells $i$ with size $A_{i}$ can occupy the same square as long as their cumulative size does not exceed its size $\sum_{i}A_{i}\leq A_{\text{max}}$. The position of a cell is not resolved beyond its grid square. Cell dynamics are defined by actions with corresponding rates, i.e., a cell proliferates with rate $\gamma$, migrates from one grid square to an adjacent one with rate $\mu$, and grows with rate $\tilde{\alpha}$.
@@
The proliferation rate $\gamma(A)$ and growth rate $\tilde{\alpha}(A)$ are both assumed to depend on the size of the cell $A$ as suggested by single-cell measurements PulHufNevStrSigFygShr2012 , while the migration rate $\mu$ is held constant. Cells interact with each other solely via exclusion principle, that is a cell can only grow or migrate if sufficient free space is available in its current grid square or the adjacent ones. This means that the values $\tilde{\alpha}$ and $\mu$ represent the maximal possible rates for growth and migration of a single cell, whereas typically these processes are mitigated or inhibited by neighboring cells competing for the same free space.
@@
As usual for a CA the cell dynamics are integrated numerically in terms of a Markov-chain-Monte-Carlo simulation with Monte-Carlo-time-step $\Delta t$. Within a time step $\Delta t$, if there are $N$ cells present, $3N$-times a random cell and a random action - proliferation, migration or growth - with corresponding rate $\omega$ are chosen. The cell then performs the chosen action with probability $\omega\Delta t$. This procedure emulates numerically that on average each cell performs each action with corresponding rate, i.e., the probability that the action has not been performed decreases exponentially $\sim\exp(-\omega t)$, given a sufficiently small time step $\Delta t\ll\text{max}\{\gamma^{-1},\mu^{-1}\}$.
I.1 Proliferation
@@
This rate increases with cell size and becomes constant at large $\gamma(A\gg A_{0})\approx\gamma_{\text{max}}$ and small sizes $\gamma(A\ll A_{0})\approx 0$, while the steepness of the transition at $A_{0}$ is set by the exponent $m$. Further, measurements of the cell areas during single-cell division (PulHufNevStrSigFygShr2012, , Inset of Fig. 4C) suggest an asymmetry between the areas of the two daughter cells, which is taken into account: If a mother cell of size $A$ divides, one of the daughter cells is assigned a fraction $f$ of the mother cell and while the other daughter is assigned the remaining part, meaning the daughter cells have size $fA$ and $(1-f)A$. Naturally, the probability density $G(f)$ of the fraction $f\in[0,1]$ is symmetric around $1/2$.
I.2 Migration
@@
I.3 Growth
Motivated by single-cell tracking measurements (PulHufNevStrSigFygShr2012, , Fig. 4A,B), we assume logistic cell growth
@@ I.3 Growth
with maximal cell size $A_{\text{max}}$ and growth rate $\alpha$. Note that the growth rate $\alpha$ indirectly affects cell proliferation, as it affects the position and width of the cell size distribution, which in turn determines the average proliferation rate according to Eq. (1). For instance, increasing the growth rate $\alpha$ leads to accelerated growth of the cell culture. While the logistic growth Eq. (3) is time-continuous and could be integrated directly via forward-Euler-scheme, we employ for simplicity the same Markov-chain-Monte-Carlo scheme as for proliferation and migration. We use $\tilde{\alpha}=1/\Delta t$ as the rate of the growth action, meaning it is performed on average once per time step and cell, and define the action as increasing the cell size by $\Delta{}A=\alpha A(1-A/A_{\text{max}})\Delta t$. Note that the time step is chosen sufficiently small such that $\mathrm{\Delta}{}A/A\ll 1$. Furthermore, the increment $\Delta{}A$ is limited to the available free space in the grid square.
II Parameter calibration

### v2

@@ Abstract
Understanding the principles of dynamic tissue organization requires an interdisciplinary effort where quantitative biological observations must be challenged by mathematical models that integrate existent insight on a theoretical level SchBru2023 ; BruArlFinRonRaeBro2021 . Cell-based mathematical models essentially consider a tissue as ’a society of cells’ Vir1855 where cells behave and interact with each other and the extracellular environment according to individual cellular characteristics summarizing the effects of sub-cellular processes, such giving rise to complex emergent processes at tissue scale. Existing cell-based model frameworks can be classified into three broad categories, (i) continuum descriptions of tissue movement which rarely describe individual cells explicitly AlaHatLowVoi2015 ; KuaGuaTanZha2023 ; Shr2005 ; BasRisJoaSasPro2009 ; HeiAleLaCZajKosCoh2020 ; KhaStr2021 , (ii) models that include an explicit description of the cell membrane, such as the cellular Potts model GraGla1992 ; LiLow2014 ; Vos2012 ; Kab2012 , vertex models NesGodWooJen2018 ; AltGanSal2017 ; FarRoeAigEatJue2007 ; KurBroZarBakFle2015 ; MorStuSabLauQueTorChaSuhRam2017 ; FleOstBakShv2014 or disk models MorStuSabLauQueTorChaSuhRam2017 ; SchMolYam2020 , and finally (iii) particle based-models or cellular automata where cells are represented as point-like particles BoeHatVosCavHerDeu2015 ; Vos2012 ; TalCavVosDeu2017 ; RosBoeLanVos2021 ; RehKliDeuVos2017 ; LanMogHueVos2022 ; HofVosRinBru2017 ; FraAlaBoeVosLan2022 ; SepPetCocGraSilHak2013 . Each class has its own merits and drawbacks: (i) Tissue scale descriptions allow to capture biophysical principles which govern large scale phenomena like tissue fluidity and deformations (elastic, plastic, viscoelastic) or jamming transitions YanBiCzaMerManMar2017 ; GniSchHal2019 . However, relating biophysical parameters of cell colonies to measurable single cell properties remains challenging and simulation of large collectives with single-cell resolution can cause high computational cost AlaHatLowVoi2015 ; KuaGuaTanZha2023 . (ii) Models that describe cells as spatially extended objects, such as cellular Potts and vertex models, can incorporate heterogeneous cell sizes and shapes within the tissue and provide a related contact network for specifying intercellular interaction. However, the cellular behavior and interaction is often described indirectly via membrane adjustments controlled by free energy terms, the latter also involving purely technical parameters which have an impact on the emergent dynamics KurBakFle2017 ; Vos2012 ; BeaKirHenGraBru2023 . This fact together with the considerable computational effort for full exploration of the parameter space limits the calibration to concrete experimental data. (iii) Cellular automata and interacting particle systems are computationally more efficient while being mathematically easier to analyze LiePalJagDra2015 . They allow to describe cellular behavior and intercellular interaction in a rule-based manner based on clear mechanistic hypotheses and with biologically meaningful parameters OsbFlePitMaiGav2017 . A major drawback, however, is the idealization of point-like cells hampering the modeling of tissue dynamics where cell sizes and shapes are supposed to play an essential role AleTre2020 .
@@
Here, we consider monolayers of growing epithelial tissue viewed as a dynamic 2D arrangement of individual cells, which grow in size, divide and migrate in response to mechanical constraints and, potentially, interaction on cell-cell contact. In-vitro experiments on the development of epithelial tissue of Madin-Darby canine kidney (MDCK) cells from Puliafito et al. (2012) PulHufNevStrSigFygShr2012 show two distinct growth regimes: At early times, in the confluent phase, cell density is low and spatially homogeneous and the colony grows exponentially in time. At longer times, in the postconfluent phase, cell density increases inside the colony and subexponential colony growth driven by more motile cells near the edge of the expanding colony is observed. The inhibition of cellular proliferation and migration in crowded environments as observed in the center of the growing monolayers is called contact inhibition of proliferation response and migration. It was found to be a not only a consequence of cell-cell contact but of mechanical constraints that cause successive cell divisions to reduce cell area. To further differentiate between the impact of biomechanical constraints and that of intercellular interaction on cell-cell contact, cell-based mathematical modeling was exploited. Puliafito et al. (2012) PulHufNevStrSigFygShr2012 used a 1D vertex model to support their findings qualitatively. However, the model parameters and the observables of the emergent behavior were not systemically related to experimental measurements. Aland et al. (2015) AlaHatLowVoi2015 could reproduce the observed cell culture topology and radial distribution function with a hybrid continuum-discrete model, however, the spatial scale of the culture had to be extrapolated to realistic culture sizes due to computational cost of the model. Jain et al. (2022) JaiWenVoi2022 incorporated proliferation in the energetic description of a multiphase field model to predict the coordination numbers of cells in growing epithelial colonies corresponding to different hypotheses on contact inhibition and to confirm the typical, experimentally observed linear boundary growth of the colony radius. The role of cell polarity and cellular cycling between motile and dividing states for contact inhibition of proliferation was explored using a disk model SchMolYam2020 . However, the question whether mechanical constraints alone, without further intercellular interaction, can explain the observed tissue-scale dynamics remains still largely open.
@@
We study this question by proposing a minimal model, an extension of probabilistic cellular automata (CA) that allows describing heterogeneous cell sizes and growth of individual cells. In our model, cells located on the nodes of a 2D regular grid can grow, divide and migrate, see Fig. 1 for an illustration. They only interact by the exclusion principle, which means that a cellular action is prevented if there is insufficient space. We calibrate the model parameters, which quantify single-cell dynamics, to previous in-vitro observations based on single-cell measurements from Puliafito et al. (2012) PulHufNevStrSigFygShr2012 , and subsequently compare the emergent features of the calibrated CA model, such as culture size and cell density, to independent experiments from the same study. These features of CA model reproduce quantitatively the corresponding experimental results, implying that the exclusion principle, the only intercellular interaction incorporated in the model, paired with size dependent proliferation rates is sufficient to generate the observed contact inhibition. We thus demonstrate the effectiveness of the CA model for efficient quantitative predictions of tissue organization on the basis of rule-based cellular behavior with heterogeneous cell sizes. We discuss the potential to incorporate effective biomechanical interactions between cells in our CA model for future studies of contact inhibition.
@@
Figure 1: Illustration of the cellular automaton (CA) model. Cells (red and green) with individual sizes $A_{i}$ are positioned on a regular grid (gray $3\times 3$ squares represent a section of this grid), where each square of the grid can be occupied by several cells as long as their total size does not exceed the size of the square $\sum_{i}A_{i}\leq A_{\text{max}}$. Each cell $i$ can proliferate with rate $\gamma(A_{i})$, migrate with rate $\mu$, and grow logistically with rate $\alpha$ to a maximal size of $A_{\text{max}}$ (each action exemplarily sketched for the green highlighted cell in the central square). Cells interact via exclusion principle, i.e., migration and growth of individual cells is inhibited by lack of sufficient free space. Dynamics are numerically integrated using Markov-chain-Monte-Carlo method with time increment $\Delta t$. Note that the position of a cell is not resolved beyond its grid square.
@@
Figure 2: Exemplary simulation from the CA model of the growth of epithelial tissue in the confluent regime (1 d, 3.3 d, 5 d), as well as at (6.2 d, 7.3 d) and post-transition (8.3 d, 8.8 d, 9.8 d) to a stationary postconfluent state. For each time point, the local cell density is encoded in color. The growing culture expands outwards over time. Around day 6 the density increases drastically in the center of the culture, as the lack of free space inhibits migration and cell growth, reflecting the transition to a postconfluent state. At the same time, a radial density gradient is visible, implying that cells at the boundary of the culture are able to grow larger and drive further outwards expansion. Model parameters calibrated as written in the main text. Note that the local average cell size is approximated by the displayed cell density as each pixel represents a lattice square of size $A_{\text{max}}=990\ \mu\text{m}^{2}$, e.g., a grid square with density of $10^{1}\cdot 10^{3}\ \text{cells/mm}^{2}$ hosts $9.9$ cells with average size $\leq$ 1/density $=100\ \mu\text{m}^{2}$.
I Data-driven approach
Our data-driven approach is based on the mentioned experiments on the development of epithelial tissue of MDCK cells PulHufNevStrSigFygShr2012 . In this study, the transition from a confluent phase to a postconfluent phase is measured during the expansion of the epithelial colony in terms of colony area, cell density, average cell cycle time $\bar{\tau}_{2}$, and average velocity of cell motion over time. Furthermore, it is observed that the cell behavior in the center of the expanding colony at the end of the confluent phase is similar to confluent, but homogeneously seeded cell cultures, except that the transition to a stationary, postconfluent state occurs soon after initiation of the culture. Thus, homogeneously seeded cultures are exploited to further quantify the long-term behavior at and after the transition from the confluent phase in terms of the asymmetry $G(f)$ of daughter cell sizes after division, distribution of cell sizes $\rho(A)$, cell size-dependent division time $\tau_{2}(A)$, and cell median size over time.
In the following, we develop an according CA model (IICA model), whose parameters are calibrated based on the measured asymmetry $G(f)$, distribution $\rho(A)$, average cell cycle time $\bar{\tau}_{2}$, cell size-dependent proliferation rate $\gamma(A)=\ln(2)/\tau_{2}(A)$, and time point of the transition $t_{T}$ (IIIParameter calibration). The emergent dynamics of the calibrated CA model are then successfully compared to the experiment in terms of colony area, cell density, and average velocity of cell motion over time for the setting of an expanding colony as well as cell median size over time for the setting of homogeneously seeded cells (IVResults). Thus, the parameter calibration (based on single-cell dynamics) and subsequent model assessment (based on emergent features of the tissue) rely on separate datasets from independent experiments.
II CA model
@@
We present an efficient, agent-based dynamical model to simulate the development of two-dimensional tissue. The model is based on a probabilistic cellular automaton (CA) but additionally allows for size changes of individual cells during cell growth and division, see Fig. 1 for an illustration of the model and Fig. 2 for an exemplary simulation of epithelial tissue. In the CA, the two-dimensional space is discretized into a regular grid of $M^{2}$ squares each with size $A_{\text{max}}$, which can be occupied by cells. Several cells $i$ with size $A_{i}$ can occupy the same square as long as their cumulative size does not exceed its size $\sum_{i}A_{i}\leq A_{\text{max}}$. While the position of a cell is not resolved beyond its grid square, the restriction of the total size ensures a monolayer without overlap of cells as in the experiment. Cell dynamics are defined by actions with corresponding rates, i.e., a cell proliferates with rate $\gamma$, migrates from one grid square to an adjacent one with rate $\mu$, and grows with rate $\tilde{\alpha}$. Further interactions could be incorporated, like cell-cell adhesion RosBoeLanVos2021 ; FraAlaBoeVosLan2022 or cell-substrate adhesion, but are neglected to focus on a minimal dynamical model.
@@
The proliferation rate $\gamma(A)$ and growth rate $\tilde{\alpha}(A)$ are both assumed to depend on the size of the cell $A$ as suggested by single-cell measurements PulHufNevStrSigFygShr2012 , while the migration rate $\mu$ is held constant. Cells interact with each other solely via exclusion principle, that is a cell can only grow or migrate if sufficient free space is available in its current grid square or the adjacent ones. This means that the values $\tilde{\alpha}(A)$ and $\mu$ represent the maximal possible rates for growth and migration of a single cell of size $A$, whereas typically these processes are mitigated or inhibited by neighboring cells competing for the same free space.
@@
The proposed model is a continuous-time cellular automaton, where updates are asynchronous with time to next reaction being an exponential random variable and its dynamics are as usually numerically approximated by a Markov-chain-Monte-Carlo simulation with Monte-Carlo-time-step $\Delta t$. Within a time step $\Delta t$, if there are $N$ cells present, $3N$-times a random cell and a random action - proliferation, migration or growth - with corresponding rate $\omega$ are chosen. The cell then performs the chosen action with probability $\omega\Delta t$. This procedure emulates numerically that on average each cell performs each action with corresponding rate, i.e., the probability that the action has not been performed decreases exponentially $\sim\exp(-\omega t)$, given a sufficiently small time step $\Delta t\ll\text{max}\{\gamma^{-1},\mu^{-1}\}$.
II.1 Proliferation
@@
This rate increases with cell size and becomes constant at large $\gamma(A\gg A_{0})\approx\gamma_{\text{max}}$ and small sizes $\gamma(A\ll A_{0})\approx 0$, while the steepness of the transition at $A_{0}$ is set by the exponent $m$. Further, measurements of the cell areas during single-cell division (PulHufNevStrSigFygShr2012, , Inset of Fig. 4C) suggest an asymmetry between the areas of the two daughter cells, which is taken into account: If a mother cell of size $A$ divides, one of the daughter cells is assigned a fraction $f$ of the mother cell and while the other daughter is assigned the remaining part, meaning the daughter cells have size $fA$ and $(1-f)A$. Naturally, the probability density $G(f)$ of the fraction $f\in[0,1]$ is symmetric around $1/2$. Note that the total cell area is not always conserved in the experiment (PulHufNevStrSigFygShr2012, , Fig. 4c, lower inset), which is ignored in the model for simplicity.
II.2 Migration
@@
II.3 Growth
We assume a generic, logistic cell growth
@@ II.3 Growth
with maximal cell size $A_{\text{max}}$ and growth rate $\alpha$. Note that the growth rate $\alpha$ indirectly affects cell proliferation, as it affects the position and width of the cell size distribution, which in turn determines the average proliferation rate according to Eq. (1). For instance, increasing the growth rate $\alpha$ leads to accelerated growth of the cell culture. While the logistic growth Eq. (3) is time-continuous and could be integrated directly via forward-Euler-scheme or the corresponding analytical solution, we employ for simplicity the same Markov-chain-Monte-Carlo scheme as for proliferation and migration. We use $\tilde{\alpha}=1/\Delta t$ as the rate of the growth action, meaning it is performed on average once per time step and cell, and define the action as increasing the cell size by $\Delta{}A=\alpha A(1-A/A_{\text{max}})\Delta t$. Note that the time step is chosen sufficiently small such that $\mathrm{\Delta}{}A/A\ll 1$. Furthermore, the increment $\Delta{}A$ is limited to the available free space in the grid square. It should be pointed out that in the previously proposed 1D vertex model, cell growth was assumed to be a consequence of the surrounding tissue stretching the cells PulHufNevStrSigFygShr2012 . In contrast, we explicitly choose a growth function that does neither rely on cells actively pushing or pulling on each other, to devise a model with minimal cell-cell interactions.
III Parameter calibration